A) A cross was made between pure breeding plants differing in two  

characters (colour and shape of seed) and the progenies were inter-crossed  

to produce F2 population. Four types of phenotypes; yellow round (1108),  

yellow wrinkled (382), green round (366) and green wrinkled (140) were  

observed in F2. Use the chi-square test to determine if these results are  

due to independent assortment.

B) Explain prophase and metaphase stages of mitosis


1
Expert's answer
2022-04-28T16:44:01-0400

A) Let's assume that the ratio of phenotypes in F2 will be 9:3:3:1.

If you count all the offsprings in F2, then there will be 1996 in total. Theoretically, the values of the number of the offspring of a particular phenotype should be as follows:

  • (1996/16)*9 = 1122.75 (against 1108);
  • (1996/16)*3 = 374.25 (against 382 and 366);
  • 1996/16 = 124.75 (against 140).


We calculate the chi-square using the formula below:

"\chi^2 = \Sigma (E-T)^2/T"


χ2 = (1108-1122.75)^2/1122.75 + (382-374.25)^2/374.25 + (366-374.25)^2/374.25 + (140-124.75)^2/124.75 = 0.19 + 0.16 + 0.18 + 1.86 = 2.39.


According to the chi-square table, with a probability greater than 0.05 and with a degree of freedom of 3, the chi-square value will be less than 7.81. 

The calculated value is less than 7.81, which means that the null hypothesis is accepted.


B) During prophase, chromosomes begin to spiralize and their functional activity decreases, as well as the shell of the nucleus is destroyed and a division spindle is formed. During metaphase, the chromosomes spiralize more strongly, attach to the spindle fibers and form the metaphase plate.